Volcán Barva (04.01.2014 to 06.30.2014)

You are here

PDF versionPDF version
Year: 
2014
Quarter: 
April - June
Communications activities: 

The Site Manager gave a talk to students of REU (Research Experience for Undergraduates) about TEAM project.

The Site Manager gave a talk about wildlife Picture Index on the International symposium of American felids Conservation.

 

The site manager elaborated a GIS layers and data in order to compile edge information for VB TEAm site. This will be a contribution for an on-going paper that is preparing Badru Mugerwa. 

The site manager attended an online training with Jimmy Maccarthy where he walked me through the features of the WPI tool and help me identify some interesting trends about the animals at our site that I can point out to my protected area manager

 

 

Unusual Events at the TEAM Site: 

excellents sightings of tapir with babies were recorded in our third sampling of terrestrial vertebrates. These animals looks very healthly which is a good indicator of the environment.

New Species: 

There are not new species for this sampling period

Protocol Activities: 

All annual census of trees, palms and lianas ≥10 cm diameter in the 1 ha plot located at  50,100, 200, 400, 600, 960, 1450, 2000 and 2800masl was completed up to date. More than 6000 stems have been measured.  A new vegetation plot located at 2400 masl was adopted. This plot was set up in 2010 following the TEAM protocol. The census done for this plot from 2010 up to date have been uploaded.

 

The climate data from Plastico (VB-CL-2) and La Selva (VB-CL-3) from April 1st  up to June 17th were collected and submitted in the new data entry form version 3.0. 

 

We have completed the terrestrial vertebrates monitoring for the sampling period in array 1, 2 an 3. (low, middle and high elevation). 60 camera traps were deployed. Interesting sightings of 18 mammals were recorded included Tapirs with babies.  3000 pictures were processed on the deskteam. All data (tpk file) were uploaded in the TEAm server.

 

 

 

Protocol Problems: 

On June we reported a damaged radiation sensor.

The sensor seems to have a bad contact. These sensors come from the factory calibrated at 200W per millivolt. The precision sensing resistor is located near the end of the wire that connect to the logger and is incorporated in a plastic molding. That means it will be hard to repair things.
The easiest would be a new sensor. As they have a fixed factory calibration the new sensor can be installed without update to the program.

 

 

Schedule Problems: 

There were not schedule problems.

Logistical Problems: 

There were not logistical problems

Suggestions: 

Initially there was an error on the climate status webpage that showed data was missing for September-December 2013 even though they were already uploaded; however, this issue appears to be resolved at this time.